# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ C.4 Proof of Lemma 11
The proof consists of three steps. In the first step, we derive a summation formula regarding the RKHS norms of $l(l\leq n)$-dimensional marginals of the process $\hat{Z}^{\tau,s}$. In the second step, we derive a decomposition for all functions in $\ell^{1}(\mathcal{X}_{n})$ using the difference formula in Lemma 10. Finally, we combine the previous two steps to finish the proof.
@@
$\bm{f}_{0}(\mathcal{X}_{l})^{T}(\hat{K}^{\tau,s}_{\mathcal{X}_{l},\mathcal{X}_{l}})^{-1}\bm{f}_{0}(\mathcal{X}_{l})=\sum_{i=1}^{l}\frac{\big{\{}f_{0}(X_{i})-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}=\bm{f}_{0}(\mathrm{pa}(X_{i}))]\big{\}}^{2}}{\mathrm{Var}\big{\{}Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big{\}}}.$ (106)
For $i=1$, note that $X_{1}$ has no parents, implying (106) directly. Next, we assume that equation (106) holds for $l=1,2,\cdots,l^{\prime}$ and we show below that it holds also for $l=l^{\prime}+1$. Define the matrix $U\in\mathbb{R}^{(l^{\prime}+1)\times(l^{\prime}+1)}$ as
@@
$(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}+1},\mathcal{X}_{l^{\prime}+1}})^{-1}$ $=U^{T}\begin{bmatrix}(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},\mathcal{X}_{l^{\prime}}})^{-1}&0\\ 0&\big{[}s^{2}-(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},X_{l^{\prime}+1}})^{T}(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},\mathcal{X}_{l^{\prime}}})^{-1}\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},X_{{l^{\prime}}+1}})\big{]}^{-1}\end{bmatrix}U.$ (107)
@@
The variance, can be rewritten along the same lines
@@
$=$ $\mathrm{Var}\big{[}Z^{\tau,s}_{X_{l^{\prime}+1}}-\mathbb{E}\big{[}{Z}^{\tau,s}_{X_{l^{\prime}+1}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{l^{\prime}+1})}\big{]}\big{]}.$ (108)
@@
Step 2 For all $0\leq j\leq\eta(n)$, we define the operator $\Theta_{j}:\ell^{1}(\mathcal{X}_{n})\to\ell^{1}(\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}})$, such that for all $f\in\ell^{1}(\mathcal{X}_{n})$,
$\Theta_{j}(f)(X_{i})=\Bigg{\{}\begin{aligned} &f(X_{i})-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}=\bm{f}(\mathrm{pa}(X_{i}))],&&X_{i}\in\mathcal{N}_{j},\\ &0,&&X_{i}\in\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}.\end{aligned}$ (109)
@@
$f=\Theta_{\eta(n)}(f)+\sum_{j=0}^{\eta(n)-1}G_{\eta(n)-1}\cdots G_{j+1}G_{j}\Theta_{j}(f).$ (110)
@@
$f|_{\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}}=\Theta_{j}(f)+\sum_{j^{\prime}=0}^{j-1}G_{j-1}\cdots G_{j^{\prime}+1}G_{j^{\prime}}\Theta_{j^{\prime}}(f),$ (111)
@@
Now suppose that (111) holds for $j=\tilde{j}$ and consider the case that $j=\tilde{j}+1$. Note that for all $X_{i}\in\cup_{j^{\prime}=0}^{\tilde{j}+1}\mathcal{N}_{j^{\prime}}$,
@@
Then in view of (111) for $j=\tilde{j}$,
@@
concluding the proof of (111) and hence also of (110).
@@
The above displays and equation (106) together with $l\leq\gamma^{\eta(l)d}$ imply,
@@
$\leq s^{-2}\gamma^{\eta(l)d}+\tau^{-2\alpha}s^{-2}\gamma^{(2\alpha-2\beta+d)\eta(l)}.$ (112)
Let $\mathcal{X}_{l}=\cup_{j=0}^{\eta(l)}\mathcal{N}_{j}$, i.e., $\mathcal{X}_{n}$ contain all elements up to layer $\eta(l)$. Let us take $f_{1}\in\mathbb{H}_{n}^{\tau,s}$ such that $f_{1}(x)=f_{0}(x),\forall x\in\mathcal{X}_{l}$ and $\Theta_{j}(f_{1})=0,\forall\eta(l)+1\leq j\leq\eta(n)$. In view of equation (110),
@@
where the last inequality utilizes Lemmas 7 and 10. Let us denote by $c$ the constant multiplier in the upper bound $\lesssim$. Then, by choosing $\eta(l)$ to be the smallest integer satisfying $c(\tau^{\alpha}\gamma^{-\eta(l)\alpha}+\gamma^{-\eta(l)\beta})\leq\epsilon$, we get
@@
If $\tau^{\alpha}\gamma^{-\eta(l)\alpha}\geq\gamma^{-\eta(l)\beta}$, then $\epsilon\asymp\tau^{\alpha}\gamma^{-\eta(l)\alpha}$, which together with (112) implies that
@@
D Proofs of the theorems and corollaries in Section 6
D.1 Classical Result for Posterior Contractions
The Bayesian posterior contraction is a well-studied problem. Here we cite two results from the literature. The first result provides sufficient conditions for Bayesian posterior contraction rates, with corresponding notations tailored for the present paper.
Lemma 15.
For arbitrary fixed $\mathcal{X}_{n}$, for $\|\cdot\|_{2,n}$ the empirical $\ell^{2}$ norm for functions defined on $\mathcal{X}_{n}$, if there exists sequences of sets $F_{n}\subset\ell^{1}(\mathcal{X}_{n})$, such that
$\ln\mathrm{Pr}(\hat{Z}_{\mathcal{X}_{n}}\not\in F_{n})\lesssim-n\epsilon_{n}^{2},$ (113)
$\ln\mathrm{Pr}(\|f_{0}-\hat{Z}_{\mathcal{X}_{n}}\|_{2,n}<2\epsilon_{n})\lesssim n\epsilon_{n}^{2},$ (114)
$\ln N(3\epsilon_{n},F_{n},\|\cdot\|_{2,n})\lesssim n\epsilon_{n}^{2},$ (115)
@@ D.1 Classical Result for Posterior Contractions
Lemma 15 is a simplified version of Theorem 8.19 of [22], with their equations (8.22) and (8.23) in place of (8.20) and (8.21), respectively, as well as noticing under the nonparametric regression settings (4), the the KL divergence of functions on $\mathcal{X}_{n}$ is equivalent to empirical $\ell^{2}$ norm on $\mathcal{X}_{n}$.
The next result is exactly Theorem 2.1 of [52] that links the three conditions of general Bayesian posterior contraction with the concentration function of Gaussian processes.
Lemma 16.
If $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\leq n\epsilon_{n}^{2}$ for some sequence $\epsilon_{n}\to 0$, then there exists sequences of subsets $F_{n}$ with $F_{n}\subset\mathbb{R}^{n}$, such that equations (113) to (115) hold.
D.2 Proof of Lemma 13
The statement is obtained by straightforward combination of Lemma 15 and 16.
D.3 Proof of Theorem 3
@@
In view of Lemma 13, it suffices to find an $\epsilon_{n}$ satisfying the inequality
$\phi^{\tau,s}_{f_{0},n}(\epsilon_{n})\lesssim n\epsilon_{n}^{2},$
Since $\phi^{\tau,s}_{f_{0},n}(\epsilon_{n})$ is upper bounded in Lemma 12, solving inequalities for each term of this upper bound yields
@@
Taking the maximum of the lower bounds on the right hand side provides the stated contraction rate. ∎
D.4 Proof of Corollary 1
@@
The case that both $\tau$ and $s$ are constants can be directly obtained from Theorem 3. For $\tau^{\alpha}s\asymp n^{\frac{\alpha-\beta}{2\beta+d}}$, by plugging it into equation (3),
@@
Therefore the posterior contraction rate is $n^{-\frac{\beta}{2\beta+d}}$. ∎
D.5 Proof of Theorem 4
We first prove a (nearly) nested property of the RKHS balls corresponding to Vecchia GPs $\hat{Z}^{\tau,s}$ in view of the scaling parameters $s$ and $\tau$. For the space scaling parameter $s$ the inclusions are strictly monotone and straightforward. However, for the time rescaling $\tau$ it is substantially more challenging. To begin with, let us denote the unit balls in the RKHS and the Banach space $(\mathcal{X}_{n},\|\cdot\|_{n,\infty})$ by
@@ D.5 Proof of Theorem 4
Lemma 17.
For arbitrary $c>0$ there exists a universal constant $C>0$, such that for all $0<\tau,\Delta\tau,s,\Delta s$ with $\Delta\tau<1$, $c<\tau-\Delta\tau\leq\tau$ and $c<s-\Delta s\leq s$,
@@
Step 1: For $f\in\mathbb{H}^{\tau,s}_{n}(1)$, in view of (35), (106) and (109),
@@
$=\sum_{i=1}^{n}\frac{[\Theta_{\eta(i)}(f)(X_{i})]^{2}}{\mathrm{Var}\big{\{}Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big{\}}}.$ (116)
Furthermore, recall the Gaussian interpolation operator $G_{j}:\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}\to\mathcal{N}_{j}$ in (33). For clarity we denote by $G^{\tau}$ the Gaussian interpolation operator used to construct $\hat{Z}^{\tau,s}$. In view of equation (110),
@@
By Lemma 7 and the assumption $\tilde{\vartheta}_{n}=O(1)$ there exists a constant $C$, such that for all $j_{1},j_{2}\in\mathbb{N}$,
$\|G^{\tau}_{j_{2}}\cdots G^{\tau}_{j_{1}}\ \|\leq\tilde{\eta}_{n},\;\;\|G^{\tilde{\tau}}_{j_{2}}\cdots G^{\tilde{\tau}}_{j_{1}}\|\leq C.$ (117)
@@
$\|G_{j}^{\tau}-G_{j}^{\tilde{\tau}}\|_{\infty}\lesssim\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}.$ (118)
Combing the previous two displays with the formulas for $f(\mathcal{X}_{n})$ and $\tilde{f}(\mathcal{X}_{n})$, we get
@@
Furthermore, in view of Cauchy-Schwarz inequality, the definition of $\Theta_{j}(f)$, assertion (D.5) and Theorem 1,
@@
$\lesssim\big{[}\Delta\tau+(\Delta\tau)^{2(\alpha-\underline{\alpha})}\big{]}s\tau^{\alpha}\ln n.$ (119)
We proceed to bound the $\mathbb{H}^{n,\tilde{\tau}}$ norm of the function $\tilde{f}$. By the definition of $\tilde{f}$, equation (D.5) and Theorem 1
@@
$\lesssim\|f\|_{\mathbb{H}^{\tau,s}_{n}}\sup_{1\leq j\leq\eta(n)}\frac{s^{2}(\tau\gamma^{-j})^{2\alpha}}{s^{2}(\tilde{\tau}\gamma^{-j})^{2\alpha}}\lesssim\|f\|_{\mathbb{H}^{\tau,s}_{n}}.$ (120)
Note that in view of (D.5) and (D.5), for all $f\in\mathbb{H}_{n}^{\tau,s}(1)$, we have $\tilde{f}\in\mathbb{H}_{n}^{\tau-\Delta\tau,s}(C)$ and $f-\tilde{f}\in\mathbb{B}_{n}\big{(}[\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}]\tau^{\alpha}s\ln n\big{)}$, for some constant $C>0$. Therefore, by writing $f=\tilde{f}+f-\tilde{f}$,
$\mathbb{H}_{n}^{\tau,s}(1)\subset\mathbb{H}_{n}^{\tau-\Delta\tau,s}(C)+\mathbb{B}_{n}\big{(}C[\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}]\tau^{\alpha}s\ln n\big{)}.$ (121)
Step 2: By the definition of the space scaling parameter $s$, (D.5) can be reformulated as
@@
$\mathbb{H}_{n}^{\tau-\Delta\tau,s}(1)\subset\mathbb{H}_{n}^{\tau-\Delta\tau,s-\Delta s}\Big{(}C\frac{1/s^{2}}{1/(s-\Delta s)^{2}}\Big{)}=\mathbb{H}_{n}^{\tau-\Delta\tau,s-\Delta s}(C^{\prime}).$ (122)
Finally, we finish the proof by combining (121) and (122). ∎
Furthermore, recall the fundamental Theorem 2.1 of [52], linking the concentration function with the three conditions of Bayesian posterior contraction rates.
We turn our attention to prove Theorem 4.
@@
$F_{n}=\cup_{(\tau,s)\in T}\mathbb{H}_{n}^{\tau,s}(a_{n})+\mathbb{B}_{n}(\epsilon_{n}),$ (123)
@@
Then, in view of Lemma 15, it suffices to prove that conditions (113)-(115) hold for $\hat{Z}^{\mathrm{Pr}}_{\mathcal{X}_{n}}$, the restriction of the process $\hat{Z}^{\mathrm{Pr}}$ to $\mathcal{X}_{n}$.
Proof of (113): In view of the definition of $\hat{Z}_{\mathcal{X}_{n}}^{\mathrm{Pr}}$ and Borell’s inequality, see Theorem 3.1 of [7],
@@
$\leq$ $\mathrm{Pr}(\tau^{\alpha}s>n^{\frac{\alpha-\beta}{2\beta+d}})+\mathrm{Pr}(\tau<n^{-\frac{\beta}{2\beta+d}}\ln^{2}n)+\mathrm{Pr}(s<1)$
$+\sup_{\tau^{\alpha}s\leq n^{\frac{\alpha-\beta}{2\beta+d}}}\Big{[}1-\Phi\Big{(}\Phi^{-1}\big{(}\mathrm{Pr}(\|\hat{Z}^{\tau,s}_{\mathcal{X}_{n}}\|_{\infty}<\epsilon_{n})\big{)}+a_{n}\Big{)}\Big{]}.$ (124)
@@
Note that minus the logarithm of the right hand side is $O(n^{\frac{d}{2\beta+d}})$.
@@
$\ln\left[\mathrm{Pr}(\tau^{\alpha}s>n^{\frac{\alpha-\beta}{2\beta+d}})+\mathrm{Pr}(\tau<n^{-\frac{\beta}{2\beta+d}}\ln^{2}n)+\mathrm{Pr}(s<1/n)\right]\lesssim-n^{\frac{d}{2\beta+d}}.$ (125)
@@
Proof of (114): Let us consider the set $\tilde{T}=\big{\{}\tau^{\alpha}s\in\big{[}n^{\frac{\alpha-\beta}{2\beta+d}},2n^{\frac{\alpha-\beta}{2\beta+d}}\big{]}\big{\}}\cap\{s\geq 1\}$. Then, in view of (38),
@@
Proof of (115): Define $\Delta\tau\in(0,1)$ such that
@@
Further define a finite set $T_{n}\subset\mathbb{R}_{+}\times\mathbb{R}_{+}$ as
$T_{n}=\left\{\Big{(}l\Delta\tau,1\Big{)}:l\in\mathbb{N},\;1\leq l\leq n^{\frac{\alpha-\beta}{(2\beta+d)\alpha}}(\Delta\tau)^{-1}\right\}.$

### v4

@@ F.4 Proof of Lemma 8
The proof consists of three steps. In the first step, we derive a summation formula regarding the RKHS norms of $l(l\leq n)$-dimensional marginals of the process $\hat{Z}^{\tau,s}$. In the second step, we obtain a decomposition for all functions in $L_{1}(\mathcal{X}_{n})$ using the difference formula in Lemma 7. Finally, we combine the results of the first two steps to complete the proof.
@@
$\bm{f}_{0}(\mathcal{X}_{l})^{T}(\hat{K}^{\tau,s}_{\mathcal{X}_{l},\mathcal{X}_{l}})^{-1}\bm{f}_{0}(\mathcal{X}_{l})=\sum_{i=1}^{l}\frac{\big\{f_{0}(X_{i})-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}=\bm{f}_{0}(\mathrm{pa}(X_{i}))]\big\}^{2}}{\mathrm{Var}\big\{Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big\}}.$ (F.14)
For $i=1$, note that $X_{1}$ has no parents, implying (F.14) directly. Next, we assume that equation (F.14) holds for $l=1,2,\cdots,l^{\prime}$ and we show below that it holds also for $l=l^{\prime}+1$. Define the matrix $U\in\mathbb{R}^{(l^{\prime}+1)\times(l^{\prime}+1)}$ as
@@
$(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}+1},\mathcal{X}_{l^{\prime}+1}})^{-1}$ $=U^{T}\begin{bmatrix}(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},\mathcal{X}_{l^{\prime}}})^{-1}&0\\ 0&\big[s^{2}-(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},X_{l^{\prime}+1}})^{T}(\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},\mathcal{X}_{l^{\prime}}})^{-1}\hat{K}^{\tau,s}_{\mathcal{X}_{l^{\prime}},X_{{l^{\prime}}+1}})\big]^{-1}\end{bmatrix}U.$ (F.15)
@@
The variance can be rewritten along the same lines
@@
$=$ $\mathrm{Var}\big[Z^{\tau,s}_{X_{l^{\prime}+1}}-\mathbb{E}\big[{Z}^{\tau,s}_{X_{l^{\prime}+1}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{l^{\prime}+1})}\big]\big].$ (F.16)
@@
Step 2 For all $0\leq j\leq\eta(n)$, we define the operator $\Theta_{j}:L_{1}(\mathcal{X}_{n})\to L_{1}(\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}})$, such that for all $f\in L_{1}(\mathcal{X}_{n})$,
$\Theta_{j}(f)(X_{i})=\Bigg\{\begin{aligned} &f(X_{i})-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}=\bm{f}(\mathrm{pa}(X_{i}))],&&X_{i}\in\mathcal{N}_{j},\\ &0,&&X_{i}\in\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}.\end{aligned}$ (F.17)
@@
$f=\Theta_{\eta(n)}(f)+\sum_{j=0}^{\eta(n)-1}G_{\eta(n)-1}\cdots G_{j+1}G_{j}\Theta_{j}(f).$ (F.18)
@@
$f|_{\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}}=\Theta_{j}(f)+\sum_{j^{\prime}=0}^{j-1}G_{j-1}\cdots G_{j^{\prime}+1}G_{j^{\prime}}\Theta_{j^{\prime}}(f),$ (F.19)
@@
Suppose that (F.19) holds for $j=\tilde{j}$ and consider the case that $j=\tilde{j}+1$. Note that for all $X_{i}\in\cup_{j^{\prime}=0}^{\tilde{j}+1}\mathcal{N}_{j^{\prime}}$,
@@
Then, in view of (F.19) for $j=\tilde{j}$,
@@
concluding the proof of (F.19) and hence also of (F.18).
@@
The above two displays together with equation (F.14) and $l\leq\gamma^{\eta(l)d}$ imply,
@@
$\lesssim s^{-2}\gamma^{\eta(l)d}+\tau^{-2\alpha}s^{-2}\gamma^{(2\alpha-2\beta+d)\eta(l)}.$ (F.20)
Let $\mathcal{X}_{l}=\cup_{j=0}^{\eta(l)}\mathcal{N}_{j}$, i.e., $\mathcal{X}_{n}$ contain all elements up to layer $\eta(l)$. Let us take $f_{1}\in\mathbb{H}_{n}^{\tau,s}$ such that $f_{1}(x)=f_{0}(x),\forall x\in\mathcal{X}_{l}$ and $\Theta_{j}(f_{1})=0,\forall\eta(l)+1\leq j\leq\eta(n)$. In view of equation (F.18),
@@
where the last inequality utilizes Lemmas 4 and 7. Let us denote by $c$ the constant multiplier in the upper bound $\lesssim$. Then, by choosing $\eta(l)$ to be the smallest integer satisfying $c(\tau^{\alpha}\gamma^{-\eta(l)\alpha}+\gamma^{-\eta(l)\beta})\leq\epsilon$, we get
@@
If $\tau^{\alpha}\gamma^{-\eta(l)\alpha}\geq\gamma^{-\eta(l)\beta}$, then $\epsilon\asymp\tau^{\alpha}\gamma^{-\eta(l)\alpha}$, which together with (F.20) implies that
@@
G Proofs of the theorems and corollaries in Section 5
G.1 Classical results for posterior contractions
Deriving contraction rates for the posterior distribution is a well-established topic. Here, we recall two fundamental results from the literature. The first provides sufficient conditions for Bayesian posterior contraction rates, with the corresponding notation adapted for the present paper.
Lemma G.1.
For arbitrary fixed $\mathcal{X}_{n}$, for $\|\cdot\|_{2,n}$ the empirical $L_{2}$ norm for functions defined on $\mathcal{X}_{n}$, if there exists a sequence of sets $F_{n}\subset L_{1}(\mathcal{X}_{n})$, such that
$\ln\mathrm{Pr}(\hat{Z}_{\mathcal{X}_{n}}\not\in F_{n})\lesssim-n\epsilon_{n}^{2},$ (G.1)
$\ln\mathrm{Pr}(\|f_{0}-\hat{Z}_{\mathcal{X}_{n}}\|_{2,n}<2\epsilon_{n})\lesssim n\epsilon_{n}^{2},$ (G.2)
$\ln N(3\epsilon_{n},F_{n},\|\cdot\|_{2,n})\lesssim n\epsilon_{n}^{2},$ (G.3)
@@ G.1 Classical results for posterior contractions
Lemma G.1 is a simplified version of Theorem 8.19 in [22], with their equations (8.22) and (8.23) replacing (8.20) and (8.21), respectively. Moreover, under the nonparametric regression settings (4), the KL divergence of functions on $\mathcal{X}_{n}$ is equivalent to empirical $L_{2}$ norm on $\mathcal{X}_{n}$.
The next result is exactly Theorem 2.1 of [57], which links the three conditions of general Bayesian posterior contraction with the concentration function of Gaussian processes.
Lemma G.2.
If $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\leq n\epsilon_{n}^{2}$ for some sequence $\epsilon_{n}\to 0$, then there exists a sequence of subsets $F_{n}\subset\mathbb{R}^{n}$, such that equations (G.1) to (G.3) hold.
G.2 Proof of Lemma 10
The statement is obtained by straightforward combination of Lemma G.1 and G.2.
G.3 Proof of Theorem 3
@@
In view of Lemma 10, it is sufficient to find an $\epsilon_{n}$ satisfying the inequality
$\phi^{\tau,s}_{f_{0},n}(\epsilon_{n})\lesssim n\epsilon_{n}^{2}.$
Since $\phi^{\tau,s}_{f_{0},n}(\epsilon_{n})$ is upper bounded in Lemma 9, solving the inequalities for each term of this upper bound yields
@@
Taking the maximum of the lower bounds on the right-hand side provides the stated contraction rate. ∎
G.4 Proof of Corollary 1
@@
The case that both $\tau$ and $s$ are constants can be directly obtained from Theorem 3. For $\tau^{\alpha}s\asymp n^{\frac{\alpha-\beta}{2\beta+d}}$, substituting this into equation (3) yields
@@
Therefore, the posterior contraction rate is $n^{-\frac{\beta}{2\beta+d}}$. ∎
G.5 Proof of Theorem 4
We first prove a (nearly) nested property of the RKHS balls corresponding to the Vecchia GPs $\hat{Z}^{\tau,s}$, with respect to the scaling parameters $s$ and $\tau$. For the space scaling parameter $s$, the inclusions are strictly monotone and straightforward. However, for the time rescaling $\tau$, the analysis is substantially more challenging. To begin with, let us denote the unit balls in the RKHS and the Banach space $(\mathcal{X}_{n},\|\cdot\|_{n,\infty})$ by
@@ G.5 Proof of Theorem 4
Lemma G.3.
For arbitrary $c>0$, there exists a universal constant $C>0$, such that for all $0<\tau,\Delta\tau,s,\Delta s$ with $\Delta\tau\leq 1$, $c\leq(\tau-\Delta\tau)/\tau$,
@@
Step 1: For $f\in\mathbb{H}^{\tau,s}_{n}(1)$, in view of (30), (F.14) and (F.17),
@@
$=\sum_{i=1}^{n}\frac{[\Theta_{\eta(i)}(f)(X_{i})]^{2}}{\mathrm{Var}\big\{Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big\}}.$ (G.4)
Furthermore, recall the Gaussian interpolation operator $G_{j}:\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}\to\mathcal{N}_{j}$ in (28). For clarity we denote by $G^{\tau}$ the Gaussian interpolation operator used to construct $\hat{Z}^{\tau,s}$. In view of equation (F.18),
@@
By Lemma 4 and the assumption $\tilde{\vartheta}_{n}=O(1)$, there exists a constant $C$, such that for all $j_{1},j_{2}\in\mathbb{N}$,
$\|G^{\tau}_{j_{2}}\cdots G^{\tau}_{j_{1}}\|\lesssim 1,\;\;\|G^{\tilde{\tau}}_{j_{2}}\cdots G^{\tilde{\tau}}_{j_{1}}\|\lesssim 1.$ (G.5)
@@
$\|G_{j}^{\tau}-G_{j}^{\tilde{\tau}}\|_{\infty}\lesssim\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}.$ (G.6)
Combing the previous two displays with the formulas for $f(\mathcal{X}_{n})$ and $\tilde{f}(\mathcal{X}_{n})$ yields
@@
Furthermore, in view of Cauchy-Schwarz inequality, the definition of $\Theta_{j}(f)$, assertion (G.5) and Theorem 1,
@@
$\lesssim\big[\Delta\tau+(\Delta\tau)^{2(\alpha-\underline{\alpha})}\big]s\tau^{\alpha}\ln n.$ (G.7)
We proceed to bound the $\mathbb{H}^{n,\tilde{\tau}}$ norm of the function $\tilde{f}$. By the definition of $\tilde{f}$, equation (G.5) and Theorem 1
@@
$\lesssim\|f\|_{\mathbb{H}^{\tau,s}_{n}}\sup_{1\leq j\leq\eta(n)}\frac{s^{2}(\tau\gamma^{-j})^{2\alpha}}{s^{2}(\tilde{\tau}\gamma^{-j})^{2\alpha}}\lesssim\|f\|_{\mathbb{H}^{\tau,s}_{n}}.$ (G.8)
Note that in view of (G.5) and (G.5), for all $f\in\mathbb{H}_{n}^{\tau,s}(1)$, we have $\tilde{f}\in\mathbb{H}_{n}^{\tau-\Delta\tau,s}(C)$ and $f-\tilde{f}\in\mathbb{B}_{n}\big([\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}]\tau^{\alpha}s\ln n\big)$, for some constant $C>0$. Therefore, by writing $f=\tilde{f}+f-\tilde{f}$,
$\mathbb{H}_{n}^{\tau,s}(1)\subset\mathbb{H}_{n}^{\tau-\Delta\tau,s}(C)+\mathbb{B}_{n}\big(C[\Delta\tau+\Delta\tau^{2(\alpha-\underline{\alpha})}]\tau^{\alpha}s\ln n\big).$ (G.9)
Step 2: By the definition of the space scaling parameter $s$, (G.5) can be reformulated as
@@
$\mathbb{H}_{n}^{\tau-\Delta\tau,s}(1)\subset\mathbb{H}_{n}^{\tau-\Delta\tau,s-\Delta s}\Big(C\frac{1/s^{2}}{1/(s-\Delta s)^{2}}\Big)=\mathbb{H}_{n}^{\tau-\Delta\tau,s-\Delta s}(C^{\prime}).$ (G.10)
Finally, we finish the proof by combining (G.9) and (G.10). ∎
Furthermore, recall the fundamental Theorem 2.1 of [57], which links the concentration function with the three conditions for Bayesian posterior contraction rates.
We turn our attention to the proof of Theorem 4.
@@
$F_{n}=\cup_{(\tau,s)\in T}\mathbb{H}_{n}^{\tau,s}(a_{n})+\mathbb{B}_{n}(\epsilon_{n}),$ (G.11)
@@
Then, in view of Lemma G.1, it suffices to prove that conditions (G.1)-(G.3) hold for $\hat{Z}^{\mathrm{Pr}}_{\mathcal{X}_{n}}$, the restriction of the process $\hat{Z}^{\mathrm{Pr}}$ to $\mathcal{X}_{n}$.
Proof of (G.1): In view of the definition of $\hat{Z}_{\mathcal{X}_{n}}^{\mathrm{Pr}}$ and Borell’s inequality, see Theorem 3.1 of [7],
@@
$\leq$ $\mathrm{Pr}(\tau^{\alpha}s>n^{\frac{\alpha-\beta}{2\beta+d}})+\mathrm{Pr}(\tau<n^{-\frac{\beta}{2\beta+d}}\ln^{2}n)+\mathrm{Pr}(s<1/n)$
$+\sup_{\tau^{\alpha}s\leq n^{\frac{\alpha-\beta}{2\beta+d}}}\Big[1-\Phi\Big(\Phi^{-1}\big(\mathrm{Pr}(\|\hat{Z}^{\tau,s}_{\mathcal{X}_{n}}\|_{\infty}<\epsilon_{n})\big)+a_{n}\Big)\Big].$ (G.12)
@@
Note that minus the logarithm of the right-hand side is $O(n^{\frac{d}{2\beta+d}})$.
@@
$\ln\left[\mathrm{Pr}(\tau^{\alpha}s>n^{\frac{\alpha-\beta}{2\beta+d}})+\mathrm{Pr}(\tau<n^{-\frac{\beta}{2\beta+d}}\ln^{2}n)+\mathrm{Pr}(s<1/n)\right]\lesssim-n^{\frac{d}{2\beta+d}}.$ (G.13)
@@
Proof of (G.2): Let us consider the set $\tilde{T}=\big\{\tau^{\alpha}s\in\big[n^{\frac{\alpha-\beta}{2\beta+d}},2n^{\frac{\alpha-\beta}{2\beta+d}}\big]\big\}\cap\{s\geq 1\}$. Then, in view of (33),
@@
Proof of (G.3): Define $\Delta\tau\in(0,1)$ such that